OpenStudy (anonymous):

Graph the following system of equations. Click on the graph until the correct one appears. x + y ≤ 1 x - y > 3

jhonyy9 (jhonyy9):

x+y<= 1 --- y<= 1-x x-y>3 ---- y<x-3 - so grafh this and will get the right answer sure

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

add values to x and see what will get for y - so like in case when x=0 result y<=1 - or in second case when x=0 result y<-3

OpenStudy (anonymous):

ok